OpenStudy (anonymous):

Help!! Describe the nature of the roots for this equation: x^2+3x-7=0

OpenStudy (anonymous):

The options are: A) Two complex roots B)Two real, irrational roots C)Two real, rational roots and D) One real, double root.

OpenStudy (anonymous):

D=3^2-4*1*-7=9+28=37>0 and not a perfect square. Hence roots are real and irrational. note:-If D>0 and a perfect square then roots are real and rational. If D=0 then roots are equal.